# On Computability of Computable Problems

### v1

@@
Abstract:Computational problems are classified into computable and uncomputable problems.If there exists an effective procedure (algorithm) to compute a problem then the problem is computable otherwise it is uncomputable.Turing machines can execute any algorithm therefore every computable problem is Turing computable.There are some variants of Turing machine that appear computationally more powerful but all these variants have been proven equally powerful.The main objective of this work is to revisit and examine the computational power of different variants of Turing machines at very fine-grain level.We achieve this objective by constructing a transform technique for Turing computable problems that transforms computable problems into another type of problems, and then we try to compute the transformed problems through different variants of Turing machine.This paper shows the existence of a realizable computational scheme that can establish a framework to analyze computational characteristics of different variants of Turing machine at infinitesimal scale.
Key Words: Turing machine, Stipulation machine, Computability, Alphabet, Tape
@@ 1 Introduction
A Turing machine consists of a finite set of states (control unit), a read-write head, and an infinite tape.A simple schematic of single tape Turing machine is shown in Fig. 1. At each step, a Turing machine reads the alphabet from tape cell that is under read-write head, changes its state, writes an alphabet on the tape cell that is under read-write head, and moves the tape head to the left $(L)$ or to the right $(R)$.If $Q$ is finite set of states and $\Gamma$ represents finite set of tape alphabets, then we can write transition function ($\delta$) of single tape Turing machine in following way.
@@
Unit time is required to execute a complete transition of transition function $\delta$.Set of input alphabet is subset of tape alphabet $\Gamma$.In the multi-tape Turing machine each tape has separate read-write head.We can write transition function of $k$-tape Turing machine in following way.
$\delta:Q\times\Gamma^{k}\longrightarrow Q\times\Gamma^{k}\times\{L,R\}^{k}$
If a problem is computable then Turing machine halts in some finite time and outputs right answer for given input.We can compute some problems more efficiently if we increase the number of tapes or alphabet size (i.e., $|\Gamma|$) [9].A single tape Turing machine can simulate $k$-tape Turing machine with quadratic slow down.If a $k$-tape Turing machine can compute a problem in $T(n)$ time then there exists a single tape Turing machine that can compute same problem in time $T(n^{2})$.For a given problem, we can achieve linear speedup (through tape compression) by increasing number of tape alphabets.If there is a problem that can be computed by some Turing machine in time $f(n)$ then, for any constant $\epsilon>0$, there exists a Turing machine with more tape alphabets that can compute the problem in time $\epsilon f(n)+n+2$.These quadratic/constant factor speedups, that we can achieve by increasing number of tapes or alphabet size, will play central role in this work.
Figure 1: Single tape Turing machine
Some variants of Turing machines appear computationally more powerful than other variants; but all these variants (defined over finite set of states) have been proven equally powerful because they can simulate each other, efficiently.This work revisits and examines the computational characteristics of different variants of Turing machines at very fine-grain level.We achieve this goal in two steps. I.Constructing a transform technique for Turing computable problems that transforms computable problems into another type of problems. II.Presenting the transformed problem to different variants of Turing machines for computation and then analysing computation of those transformed problems.The scope of this work is being restricted and it includes only deterministic variants Turing machines.
@@ 2 Construction of Transform
Suppose there is a deterministic single tape Turing machine $D_{R}$ and it is defined over binary input.The unit time that $D_{R}$ takes to complete one transition of its transition function is represented by $T_{t}$, and the time $T_{t}$ is constant and same for every Turing machine mentioned in this paper.Turing machine $D_{R}$ computes a partially computable function $R$ that is not total.For a given input, the Turing machine $D_{R}$ requires $N_{R}$ number of steps to compute $R$.If $R$ is computable for given input then $N_{R}$ is some finite number, and it depends on input length.For a given input, the function $R$ can be either computable or uncomputable.Total time required to compute $R$ for a given input is represented by $f(R)$.
@@
With a given input, there are two possible outcomes for function $R$, either $R$ is a computable function (if $D_{R}$ computes and halts) or $R$ is uncomputable function (if $D_{R}$ loops).We can state these two possibilities in term of $f(R)$ in following way.For a given input
@@
    If $R$ is computable then $f(R)$ has finite value
@@
    If $f(R)$ has infinite value (i.e., $N_{R}\rightarrow\infty$ ) then $R$ is uncomputable
Suppose, function $P$ in (2) is a computable (partially computable and total) function.If $P$ is computable then $f(P)$ has some finite value for every input.
@@
Suppose, we have a variable $F$ such that $f(P,F)=N_{P}T_{t}F_{P}$ and $1\leq F_{P}<\infty$.If $M_{P}=\nicefrac{{1}}{{F_{P}}}$ then $f(P,M)=\nicefrac{{N_{P}T_{t}}}{{M_{P}}}$ and $0<M_{P}\leq 1$.This variable $M$ will be explained as we move forward in construction of this transform.The behaviour of (3), when $M_{P}=1$ and $M_{P}\rightarrow 0$, is represented in (4) and (5), respectively.
@@
Suppose, the new function $f(P,M)$ represents time required to compute some partially computable function $P(m)$ for a given input.Given that $P$ is computable function and $P(m)$ is partially computable function, we can summarise (4 and 5) in following way.
For a given input
@@
In (3), if $M_{P}=1-2^{-X_{P}}$ and $0<M_{P}\leq 1$ then $X_{P}=-\log_{2}(1-W)$ and $0<X_{p}\leq\infty$. After replacing $M_{P}$ with $1-2^{-X_{P}}$ in (3) we get (6).Given that $P$ is computable function and $P(x)$ is partially computable function.
@@
If it is given that $P$ is computable function and $P(x)$ is partially computable function then we can summarise (7 and 8) in following way.
For a given input
@@
We can implement partially computable and computable (partially computable and total) functions through Turing machines.For a given input, there always exists a Turing machine $M_{1}$ that can execute (a function or a decision) task $Y$ if there exists a partial computable function for $Y$ and $M_{1}$ halts on the given input.Now we present Stipulation machine that is also a realizable computational object, just like Turing machines; and Turing machines interacts with Stipulation machines through some mutually binding set of rules (postulates).We can materialize the mathematical formulations presented in (6,7 and 8) through a realizable computational scheme that consists of Turing machine, Stipulation machine and some postulates.
@@ 2.1 Stipulation machine
There is a Turing machine $M_{1}$ that computes problem $Y$, and it is connected to Stipulation machine.A simple schematic of interconnection between Turing machine and Stipulation machine is shown in Fig.2. There are eight postulates that describe the interactions between Turing machine and Stipulation machine. I. Stipulation machine writes input on the space specified for input in Turing machine $M_{1}$, and there is one and only space for input in $M_{1}$. II. Stipulation machine writes only those inputs that $M_{1}$ can decide, i.e., $M_{1}$ is total Turing machine for the Stipulation machine. III.Stipulation machine rewrites input in Turing machine $M_{1}$ after every fixed time interval $T_{Y}$. IV.Length of every recurring input is equal to input length of first input of a computation. V. No two consecutive inputs of Turing machine $M_{1}$ can be identical in a computation. VI. When Turing machine $M_{1}$ halts, it informs Stipulation machine. VII. When Stipulation machine receives halting signal, it stops rewriting input on Turing machine $M_{1}$ for the computation. VIII. We can write input in $M_{1}$ only through Stipulation machine when it is connected to Stipulation machine.
The time $T_{Y}$ is characteristic of Stipulation machine and $0<T_{Y}\leq\infty$, and $T_{Y}=\infty$ means Stipulation machine does not rewrite input.If there is a Turing machine $M_{1}$ that is connected with Stipulation machine then it is represented by $M_{1}(t)$, and $\langle M_{1}(t)\rangle=\langle M_{1}\rangle$.Just like Turing machine, Stipulation machine is also a realizable object.Therefore, we can construct the computational scheme that is based on the eight postulates and interactions presented in Fig. 2, and we can observe the conclusions that are based this computational scheme.
@@
2.2 Computation by a Turing machine $M(t)$
There is a Turing machine $P_{r}$ that accepts input if sum of input binary string is even and it rejects if sum of input binary strings is odd.Turing machine $P_{r}$ is a decider if input is $a$ for all $a\in A$.
$A=\{1101+101,11+10111,101+1101,1+11+1011,110+111+100,\ldots\}$
We give an input string $a$ (e.g., $11+10111$) to $P_{r}$.The Turing machine $P_{r}$ decides and halts when given input is $a$, i.e., $P_{r}(a)$ is decidable.Now, we transform $P_{r}$ into $P_{r}(t)$ by connected it to Stipulation machine and give input $a$ through Stipulation machine. In case of $P_{r}(t)$ with input $a$, the behaviour of the decidability is quite different when $T_{P_{r}}$ is relatively large and $T_{P_{r}}$ is relatively small.In following two cases we analyse the decidability of $P_{r}(t)$ with input $a$ and $0<T_{P_{r}}\leq\infty$; particularly, when $T_{P_{r}}\to 0$ and $T_{P_{r}}=\infty$.
@@ 2.2 Computation by a Turing machine $M(t)$
Stipulation machine writes input string $a$ on Turing machine $P_{r}(t)$, the Turing machine $P_{r}(t)$ decides and halts on input $a$ in time less than $T_{P_{r}}$, Stipulation machines receives halts signal from $P_{r}(t)$ and does not rewrite input.If $T_{P_{r}}$ is relatively large then there is no difference between $P_{r}$ and $P_{r}(t)$, i.e., both $P_{r}(a)$ and $P_{r}(t)(a)$ are decidable.
$L_{1}=\{\langle P_{r}(t),a\rangle|\text{ Turing machine }P_{r}(t)\text{ decides }a\}$
With input $a$ and relatively large $T_{P_{r}}$ the language $L_{1}$ decidable; both, $P_{r}(a)$ and $P_{r}(t)(a)$ are decidable.There is no difference between $P_{r}$ and $P_{r}(t)$ when $T_{P_{r}}=\infty$.Therefore, $P_{r}(t)=P_{r}$ when $T_{P_{r}}=\infty$.
@@
If $T_{P_{r}}$ is relatively too small then inputs of $P_{r}(t)$ change too quickly.Now, the time $T_{P_{r}}$ is relatively so small that Stipulation machine rewrites input on $P_{r}(t)$ far before $P_{r}(t)$ decides and halts on an input.Consequently, $P_{r}(t)$ carries out computation on some parts of each recurring input in some order.The transitions that $P_{r}(t)$ takes depends on its transition function, current configuration and those some parts of recurring inputs.Those some parts and some order make arbitrary string $f$ of some arbitrary length, and composition of $f$ depends on $T_{P_{r}}$.The string $f$ may not be element of set $A$.
@@
Only end points of $T_{P_{r}}$ (i.e., $T_{P_{r}}\to 0$ and $T_{P_{r}}=\infty$) are considered in following subsection but $T_{P_{r}}$ can have any value on interval $(0\quad\infty]$ and it is discussed in next section (Section 3).The time $P_{r}$ takes to recognize $a$ in $P_{r}(a)$ is represented by $f(P_{r})$, and the time $P_{r}(t)$ takes to recognize $a$ in $P_{r}(t)(a)$ is represented by $f(P_{r},T)$.Following mathematical expression can represent the conclusions of case I and II.
@@
Stipulation machine and Turing machine are realizable computational objects.Therefore, Stipulation machines and Turing machines can be designed to replicate these phenomena and conclusions.Now we describe some examples where there is some problem $P$ that is computable for input $\psi$, and we transform $P(\psi)$ into $P(t)(\psi)$.The problem $P$ can be a decision or a function problem.
@@ 2.3 Examples of Transformed Problems
The problem $P_{1}$ is to twice evaluate the satisfiability of a given formula (e.g., $\Phi$ in 10) for a given input ($n_{\Phi}$), if both answers are yes then accept, if both answers are no then reject; and if both answers are no same then start evaluating again and continue until we have two consecutive same answers.If $P_{1}$ is decidable with given input ($\Phi$ and $n_{\Phi}$) then it will remain decidable if we evaluate the satisfiability for any finite number of times instead of two times, because we will have same answers (yes/no) irrespective how many time we evaluate given formula for the given input. We transform $P_{1}$ into $P_{1}(t)$, and variation time of Stipulation machine for $P_{1}(t)$ is represented by $T_{P_{1}}$.We give $\Phi$ and $n_{\Phi}$ for which the problem $P_{1}$ is decidable, and then through stipulation machine we give same input ($\Phi$ and $n_{\Phi}$) to $P_{1}(t)$
@@
Given that ${P_{1}}(\Phi,n_{\Phi})$ is decidable
@@
    If $T_{P_{1}}=\infty$ then $f(P_{1},T)=f(P_{1})$; consequently, $P_{1}(t)(\Phi,n_{\Phi})$ is decidable problem
@@
    If $T_{P_{1}}\rightarrow 0$ then $f(P_{1},T)\to\infty$; consequently, $P_{1}(t)(\Phi,n_{\Phi})$ is undecidable problem
Through this transform technique, we can map a decidable decision problem into a decidable problem or an undecidable problem.Suppose there are three $n\times n$ non-zero matrices $A$, $B$, and $C$ such that $AB=C$.The problem $P_{2}$ is to compute $C$ four times for given $A$ and $B$, if in all four answers $(c_{11}\geq c_{nn})\land(c_{1n}<c_{n1})$ then accept, if in all four answers $(c_{11}<c_{nn})\land(c_{1n}\geq c_{n1})$ then reject; and if all four answers are not same then compute $C$ again from start and continue until we have four consecutive same answers.We transform $P_{2}$ into $P_{2}(t)$, and variation time of Stipulation machine for $P_{2}(t)$ is represented by $T_{P_{2}}$. We give $A$ and $B$ for which $P_{2}(A,B)$ is decidable, and then we gave same input $(A,B)$ to $P_{2}(t)$ through Stipulation machine.
{paracol}
2
[…]
@@
Suppose there are two $n\times n$ matrices $A$ and $B$, and both matrices have integer entries.The problem $P_{3}$ is to compute $AB$ thrice, and in all three computations, if $AB=I$ then accept, if $AB\neq I$ then reject; and if all three answers are not same then compute $AB$ again from start and continue until we have three consecutive same answers.We transform $P_{3}$ into $P_{3}(t)$, and variation time of Stipulation machine for $P_{3}(t)$ is represented by $T_{P_{3}}$. We give $A$ and $B$ for which $P_{3}(A,B)$ is decidable, and then we gave same input to $P_{3}(t)$ through Stipulation machine. {paracol}2
@@
So far, we have applied this transform technique on decidable decision problems, but we can also apply it on computable functions problems.Suppose there is a set $u$ that consists of finite number of finite length binary strings, and each string has equal number of ones.The problem $P_{4}$ is to compute the number of ones in each string, if number of ones are equal in each string then output the result and halt, and if number of ones in each string are not equal then start computing again and continue until equal number of ones computed in each string.We transform $P_{4}$ into $P_{4}(t)$, and variation time of Stipulation machine for $P_{4}(t)$ is represented by $T_{P_{4}}$.The problem $P_{4}$ is computable for input $u$, and we give input $u$ to $P_{4}(t)$ through Stipulation machine.
@@
Suppose there is an $m\times n$ matrix A with $a_{ij}\in\mathbb{N}$.The problem $P_{5}$ is to compute $det(A)$ through Gaussian elimination, and it is a computable problem when input is $A$.We transform $P_{5}$ into $P_{5}(t)$, and variation time of Stipulation machine for $P_{5}(t)$ is represented by $T_{P_{5}}$.We give $A$ for which $P_{5}$ is computable, and then we give $A$ to $P_{5}(t)$ through Stipulation machine.
$|A|=\left|\begin{array}[]{cccccc}a_{11}&a_{12}&a_{13}&\cdots&a_{1n}\\ a_{21}&a_{22}&a_{23}&\cdots&a_{2n}\\ a_{31}&a_{32}&a_{33}&\cdots&a_{3n}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{m1}&a_{m2}&a_{m3}&\cdots&a_{mn}\end{array}\right|$ (26)
$|A_{d}|=\left|\begin{array}[]{cccccc}a_{11}(t)&a_{12}(t)&a_{13}(t)&\cdots&a_{1n}(t)\\ a_{21}(t)&a_{22}(t)&a_{23}(t)&\cdots&a_{2n}(t)\\ a_{31}(t)&a_{32}(t)&a_{33}(t)&\cdots&a_{3n}(t)\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{m1}(t)&a_{m2}(t)&a_{m3}(t)&\cdots&a_{mn}(t)\end{array}\right|$ (27)
@@
Suppose there is an $n\times n$ matrix $Q$ with $q_{ij}\in\mathbb{N}$, and rank of $Q$ is $n$.The problem $P_{5}$ is to compute inverse of a given matrix through Gaussian elimination, and $P_{5}$ is computable when input is $Q$.Because $Q$ is invertible matrix and $(Q^{-1})^{-1}=Q$ therefore $P_{6}(P_{6}(Q))=Q$. If $A$ and $B$ are two invertible matrices and $A\neq B$ then $P_{6}(A)\neq P_{6}(B)$.We transform $P_{6}$ into $P_{6}(t)$, and variation time of Stipulation machine for $P_{6}(t)$ is represented by $T_{P_{6}}$.We give $Q$ for which $P_{6}$ is computable, and then we give $Q$ to $P_{6}(t)$ through Stipulation machine.
@@
Suppose there is a problem that has partially computable function $P$, and $P$ is computable when input is $\psi$. We transform $P(\psi)$ into $P(t)(\psi)$ through a transform that will be represented by $f(Z,T)$; and transformed problem $P(t)(\psi)$ can be computable or uncomputable, depending upon time $T_{P}$.

### v3

@@
On Computability of Computable Problems
Asad Khaliq
E-mail address: asadmalik@giki.edu.pk
Faculty of Electrical Engineering, GIK Institute of Engineering Sciences and Technology. KPK, Topi, 23640, Pakistan.
{adjustwidth}
50pt50pt Abstract: Computational problems are classified into computable and uncomputable problems.If there exists an effective procedure (algorithm) to compute a problem then the problem is computable otherwise it is uncomputable.Turing machines can execute any algorithm therefore every computable problem is Turing computable.Cardinality of Turing machines and computable problems is equal-both are countably infinite.In this paper we introduce new type of problems by constructing a transform technique and applying it on some computable problems.The transformed problems can be computable of uncomputable.
Key Words: Turing machine,Transform, Stipulation machine, Computability
@@ 1 Introduction
A Turing machine consists of a finite set of states (control unit), a read-write head, and an infinite tape.A simple schematic of single tape Turing machine is shown in Fig. 1. At each step a Turing machine reads the alphabet from tape cell that is under read-write head, changes its state, writes an alphabet on the tape cell that is under read-write head, and moves the tape head to the left $(L)$ or to the right $(R)$.If $Q$ is finite set of states and $\Gamma$ represents finite set of tape alphabets then we can write transition function ($\delta$) of single tape Turing machine in following way.
@@
Unit time is required to execute one complete transition of transition function $\delta$.This unit time is same for every single transition of a transition function.If a problem is computable for given input then Turing machine halts in some finite time and outputs right answer.Now we construct a transform technique and apply it on some computable problems, and then we explain computability of the transformed problems.
@@ 2 Construction of Transform
Suppose there is a deterministic Turing machine $D_{R}$.The unit time that $D_{R}$ takes to complete one transition of its transition function is represented by $T_{t}$, and the time $T_{t}$ is constant and same for every Turing machine mentioned in this paper.Turing machine $D_{R}$ computes a partially computable function $R$.Suppose, for a given input $y$, the Turing machine $D_{R}$ requires $N_{R}$ number of steps to compute $R(y)$.If $R(y)$ is computable then $N_{R}$ is some finite number.For input $y$, the partially computable function $R$ can be either computable or uncomputable.Total time required to compute $R(y)$ by $D_{R}$ for input $y$ is represented by $f(R)$.
@@
With input $y$, there are two possible outcomes for partially computable function $R$, either $R(y)$ is a computable (if $D_{R}$ computes $R(y)$ and halts) or $R(y)$ is uncomputable (if $D_{R}$ loops forever).We can state these two possibilities in term of $f(R)$ in following way.For given input $y$
@@
    If $R(y)$ is computable then $f(R)$ has finite value
@@
    If $f(R)$ has infinite value (i.e., $N_{R}\rightarrow\infty$ ) then $R(y)$ is uncomputable
There is another a function $P$ and it is a total computable.If $P$ is total computable function then $f(P)$ has some finite value for every input.
@@
Suppose we have a variable $F$ such that $f(P,F)=N_{P}T_{t}F_{P}$ and $1\leq F_{P}<\infty$.If $M_{P}=\nicefrac{{1}}{{F_{P}}}$ then $f(P,M)=\nicefrac{{N_{P}T_{t}}}{{M_{P}}}$ and $0<M_{P}\leq 1$.This variable $M$ will be explained as we move forward in construction of this transform.The behaviour of (3), when $M_{P}=1$ and $M_{P}\rightarrow 0$, is represented in (4) and (5), respectively.
@@
The new function $f(P,M)$ represents time required to compute function $P(m)$ for given input and $M$.Given that $P$ is computable function and $P(m)$ is new function, we can summarise (4) and (5) in following way.
For given input
@@
In (3), if $M_{P}=1-2^{-X_{P}}$ and $0<M_{P}\leq 1$ then $X_{P}=-\log_{2}(1-M_{P})$ and $0<X_{p}\leq\infty$. After replacing $M_{P}$ with $1-2^{-X_{P}}$ in (3) we get (6).For any given input, $P$ is computable and computability of $P(x)$ depends on $x$, and in computation they require time $f(P)$ and $f(P,X)$, respectively.
@@
If it is given that $P$ is computable function but computability of $P(x)$ depends on value of $x$ then we can summarise (7 and 8) in following way.
For given input
@@
We can implement partially computable and computable (partially computable and total) functions through Turing machines.For a given input, there always exists a Turing machine $M_{1}$ that can compute task $Y$ if there exists a computable function for $Y$.We can materialize the mathematical formulations presented in (6,7 and 8) through a realizable computational scheme that consists of Stipulation machine and Turing machine.Stipulation machine is a realizable computational object, just like Turing machines.Stipulation machine interacts with a Turing machine through some mutually binding set of rules (postulates).
@@ 2.1 Stipulation machine
Suppose, there is a Turing machine $M_{1}$ that computes some problem $Y$, and $M_{1}$ is connected to Stipulation machine.A simple schematic of interconnection between Turing machine and Stipulation machine is shown in Fig.2. There are eight postulates that describe the interactions between Turing machine and Stipulation machine. I. Stipulation machine writes input on the space specified for input in Turing machine $M_{1}$, and there is one and only space for input in $M_{1}$. II. Stipulation machine writes only those inputs that $M_{1}$ can decide, i.e., $M_{1}$ is total Turing machine for the Stipulation machine. III.Stipulation machine rewrites input in Turing machine $M_{1}$ after every fixed time interval $T_{Y}$. IV.Length of every recurring input is equal to input length of first input in a computation. V. No two consecutive inputs of Turing machine $M_{1}$ can be identical in a computation. VI. When Turing machine $M_{1}$ halts the Stipulation machine gets this information. VII. When Stipulation machine receives halting signal, it stops rewriting input on Turing machine $M_{1}$ for current computation. VIII. We can write input in $M_{1}$ only through Stipulation machine when it is connected to Stipulation machine.
The time $T_{Y}$ is characteristic of Stipulation machine and $0<T_{Y}\leq\infty$, and $T_{Y}=\infty$ means Stipulation machine does not rewrite input in a computation.If there is a Turing machine $M_{1}$ that is connected with Stipulation machine then it is represented by $M_{1}(t)$, and $\langle M_{1}(t)\rangle=\langle M_{1}\rangle$.Just like Turing machines, Stipulation machines is also a realizable object.Therefore, we can construct the computational scheme that is based on the eight postulates and interactions presented in Fig.2.It is important to note that Turing machine has only one place for input when it is connected to Stipulation machine and that place is accessible to Stipulation machine.
@@
2.2 Computation by a Turing machine $P_{r}$ and $P_{r}(t)$
There is a Turing machine $P_{r}$ that computes sum of input binary strings twice, it accepts if sums are even and equal in both computations, it rejects if sums are odd and equal in both computations, and it computes again from the start if sums in both computations is not identical (even and equal or odd and equal).Suppose there is set $A$ and $a\in A$, and $P_{r}$ decides for every input $a$.
$A=\{110+101010,1100+100,1010+110,100+10,1101110+10,\ldots\}$
We give an input string $a$ (e.g., $1100+10110$) to $P_{r}$.The Turing machine $P_{r}$ decides and halts when given input is from $A$, i.e., $P_{r}(a)$ is decidable.Now, we transform $P_{r}$ into $P_{r}(t)$ by connected it to Stipulation machine and give input $a$ through Stipulation machine.In case of $P_{r}(t)$ with input from $A$, the behaviour of the decidability is quite different when $T_{P_{r}}$ is relatively large and $T_{P_{r}}$ is relatively small.In following two cases we analyse the decidability of $P_{r}(t)$ with input from $A$ and $0<T_{P_{r}}\leq\infty$; particularly, when $T_{P_{r}}\to 0$ and $T_{P_{r}}=\infty$.
@@ 2.2 Computation by a Turing machine $P_{r}$ and $P_{r}(t)$
Stipulation machine writes input string $a$ on Turing machine $P_{r}(t)$, the Turing machine $P_{r}(t)$ decides and halts on input $a$ in time less than $T_{P_{r}}$, Stipulation machines receives halts signal from $P_{r}(t)$ and does not rewrite input.If $T_{P_{r}}$ is relatively large then there is no difference between $P_{r}$ and $P_{r}(t)$, i.e., both $P_{r}(a)$ and $P_{r}(t)(a)$ decide and halt.
With $T_{P_{r}}=\infty$ and input $a$, for all $a\in A$, both $P_{r}(a)$ and $P_{r}(t)(a)$ are decidable.There is no difference between $P_{r}$ and $P_{r}(t)$ when $T_{P_{r}}=\infty$.Therefore, $P_{r}(t)=P_{r}$ when $T_{P_{r}}=\infty$.
@@
If $T_{P_{r}}$ is relatively too small then inputs of $P_{r}(t)$ change too quickly.Now, the time $T_{P_{r}}$ is relatively so small that Stipulation machine rewrites input on $P_{r}(t)$ far before $P_{r}(t)$ decides and halts on an input.Consequently, $P_{r}(t)$ carries out computation on some parts of each recurring input in some order.The transitions that $P_{r}(t)$ takes depends on its transition function, current configuration and those some parts of recurring inputs.Those some parts and some order make $P_{r}(t)$ to take transitions that it would take if input is some unknown string $f$, and length and composition of $f$ depends on $T_{P_{r}}$.The arbitrary string $f$ may not be element of set $A$.
@@
Only end points of $T_{P_{r}}$ (i.e., $T_{P_{r}}\to 0$ and $T_{P_{r}}=\infty$) are considered here but $T_{P_{r}}$ can have any value on interval $(0\quad\infty]$ and it is discussed in next section (Section 3).The time $P_{r}$ takes to decide $a$ in $P_{r}(a)$ is represented by $f(P_{r})$, and the time $P_{r}(t)$ takes to recognize $a$ in $P_{r}(t)(a)$ is represented by $f(P_{r},T)$.Following mathematical expression can represent the conclusions of case I and II.
@@
Stipulation machine and Turing machine are realizable computational objects.Therefore, Stipulation machines and Turing machines can be designed to experimentally validate this transform.Now we construct and describe some examples where a problem $P(\psi)$ is computable for input $\psi$, and we transform $P(\psi)$ into $P(t)(\psi)$.The problem $P$ can be a decision or a function problem.
@@ 2.3 Examples of Transformed Problems
The problem $P_{1}$ is to evaluate a 3-SAT (e.g., 12) twice for a given input $n_{\Phi}$, if both answers are yes then accept, if both answers are no then reject; and if both answers are not same then start evaluating again and continue until we have two consecutive same answers.If $P_{1}$ is decidable for given input $n_{\Phi}$ then it will remain decidable if we compute it for any number of times instead of two times-because we will have same answer (yes or no) irrespective how many time we evaluate a formula on a given input.We transform $P_{1}$ into $P_{1}(t)$, and variation time of Stipulation machine for $P_{1}(t)$ is represented by $T_{P_{1}}$.In computation of $P_{1}(t)(n_{\Phi})$, we give input $n_{\Phi}$ to the Turing machine through stipulation machine.
@@
Given that ${P_{1}}(n_{\Phi})$ is decidable
@@
    If $T_{P_{1}}=\infty$ then $f(P_{1},T)=f(P_{1})$; consequently, $P_{1}(t)(n_{\Phi})$ is decidable problem
@@
    If $T_{P_{1}}\rightarrow 0$ then $f(P_{1},T)\to\infty$; consequently, $P_{1}(t)(n_{\Phi})$ is undecidable problem
Even when $P_{1}(t)(n_{\Phi})$ is decidable for every recurring input $n_{\Phi}$ but $P_{1}(t)$ may never halt if $T_{P_{1}}$ is too small.Through this transform technique, we can map a decidable decision problem into a decidable problem or an undecidable problem.Suppose there are two $n\times n$ non-zero matrices $A$, and $B$ such that $AB=C$.The problem $P_{2}$ is to compute $C$ four times for given $A$ and $B$, if in all four answers $c_{11}c_{1n}\geq c_{nn}c_{n1}$ then accept, if in all four answers $c_{11}c_{1n}<c_{nn}c_{n1}$ then reject; and if all four answers are not same then compute $C$ again from start and continue until we have four consecutive same answers.We transform $P_{2}$ into $P_{2}(t)$, and variation time of Stipulation machine for $P_{2}(t)$ is represented by $T_{P_{2}}$.In computation of $P_{2}(t)(A,B)$, we give input ($A$,$B$) to the Turing machine through stipulation machine.Decidability of transformed problem $P_{2}(t)(A,B)$ depends on $T_{P_{2}}$.
@@
Suppose there are two $n\times n$ matrices $A$ and $B$, and both matrices have integer entries.The problem $P_{3}$ is to compute $AB$ thrice, and in all three computations, if $AB=I$ then accept, if $AB\neq I$ then reject; and if all three answers are not same then compute $AB$ again from start and continue until we have three consecutive same answers.We transform $P_{3}$ into $P_{3}(t)$, and variation time of Stipulation machine for $P_{3}(t)$ is represented by $T_{P_{3}}$.In computation of $P_{3}(t)(A,B)$, we give input ($A$, $B$) to the Turing machine through stipulation machine.Decidability of $P_{3}(t)(A,B)$ depends on variation time $T_{P_{3}}$. {paracol}2
@@
So far, we have applied this transform technique on decidable decision problems, but we can also apply it on computable functions problems.Suppose there is some binary string $u$.The problem $P_{4}$ is to compute the number of ones in $u$ for $K$ times for some $K\in\mathbb{N}$, if number of ones are equal in each count then output the result and halt, and if number of ones in each count are not equal then start counting again and continue until we have equal number of ones in all $K$ counts and output the count.We transform $P_{4}$ into $P_{4}(t)$, and variation time of Stipulation machine for $P_{4}(t)$ is represented by $T_{P_{4}}$.Given that the problem $P_{4}(u)$ is computable, and we give input $u$ to $P_{4}(t)$ through Stipulation machine for computation of $P_{4}(t)(u)$.Just like in any other transformed problem the value of time $T_{P_{4}}$ plays decisive role in computability of $P_{4}(t)(u)$.
@@
Suppose there is an $m\times n$ matrix A with $a_{ij}\in\mathbb{N}$, and A is invertible matrix.The problem $P_{5}$ is to compute $A^{-1}$, prove $AA^{-1}=I$ and halt, and if $AA^{-1}\neq I$ then start again and continue until $AA^{-1}=I$ for some invertible matrix $A$.The problem $P_{5}$ is a computable when input is $A$.We transform $P_{5}$ into $P_{5}(t)$, and variation time of Stipulation machine for $P_{5}(t)$ is represented by $T_{P_{5}}$.For input $A$ the problem $P_{5}(A)$ is computable, and we give $A$ to $P_{5}(t)$ through Stipulation machine in computation of $P_{5}(t)(A)$.
$A=\left[\begin{array}[]{cccccc}a_{11}&a_{12}&a_{13}&\cdots&a_{1n}\\ a_{21}&a_{22}&a_{23}&\cdots&a_{2n}\\ a_{31}&a_{32}&a_{33}&\cdots&a_{3n}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{m1}&a_{m2}&a_{m3}&\cdots&a_{mn}\end{array}\right]$ (26)
$A(t)=\left[\begin{array}[]{cccccc}a_{11}(t)&a_{12}(t)&a_{13}(t)&\cdots&a_{1n}(t)\\ a_{21}(t)&a_{22}(t)&a_{23}(t)&\cdots&a_{2n}(t)\\ a_{31}(t)&a_{32}(t)&a_{33}(t)&\cdots&a_{3n}(t)\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{m1}(t)&a_{m2}(t)&a_{m3}(t)&\cdots&a_{mn}(t)\end{array}\right]$ (27)
$f(P_{5},T)=\frac{N_{P_{5}}T_{t}}{1-2^{-T_{P_{5}}}}$ (28)
$\lim_{{T_{P_{5}}=\infty}}f(P_{5},T)=f(P_{5})$ (29)
$\lim_{{T_{P_{5}}\to 0}}f(P_{5},T)\to\infty$ (30)
@@
Suppose there is an $n\times n$ matrix $Q$ with $q_{ij}\in\mathbb{N}$, and rank of $Q$ is $n$.The problem $P_{6}$ is to compute inverse of input matrix and then inverse of inverse matrix, if double inverse matrix is identical to input matrix then halt otherwise compute this task again from start on input matrix.The problem $P_{6}$ is computable if input is $Q$ because $(Q^{-1})^{-1}=Q$.We transform $P_{6}$ into $P_{6}(t)$, and variation time of Stipulation machine for $P_{6}(t)$ is represented by $T_{P_{6}}$.There is $Q$ for which $P_{6}(Q)$ is computable, and then we give $Q$ to $P_{6}(t)$ through Stipulation machine for computation of $P_{6}(t)(Q)$.
@@
Suppose there is a problem $P(\psi)$ and it is computable with input $\psi$.We transform this computable problem $P(\psi)$ into $P(t)(\psi)$ through a transform that will be represented by $f(Z,T)$; and transformed problem $P(t)(\psi)$ can be computable or uncomputable, depending on time $T_{P}$.